OpenStudy (anonymous):

Find the sum of the following series: \[1 - \frac{\ln \pi}{1!} + \frac{\ln^{2}\pi}{2!} - \frac{\ln^{3}\pi}{3!} + ...\]

OpenStudy (vishal_kothari):

from did u get this question bcoz same was in purdue?

OpenStudy (anonymous):

i got it as a bonus question from my exam earlier.

OpenStudy (vishal_kothari):

then u must be in college....

OpenStudy (anonymous):

well... yeah

OpenStudy (anonymous):

this sereis converges, bt how to find the sum?

OpenStudy (phi):

See if you can make progress with definition 2 http://en.wikipedia.org/wiki/Characterizations_of_the_exponential_function

OpenStudy (anonymous):

then the answer is e^(ln pi) right?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

e^(-ln pi) is correrct?

OpenStudy (anonymous):

which is?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Final Answer 0.3183098862..

OpenStudy (anonymous):

\[e^{-\ln(\pi)}=\frac{1}{\pi}\] maybe

OpenStudy (anonymous):

yep \[0.3183098862.. = \frac{1}{\pi}\]

OpenStudy (anonymous):

Then I made it..

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

you have any other ques. like this?